Jessica's grandmother and Jessica have the same birthday. Jessica's grandmother was 54 when Jessica was born. In exactly 1 year, grandmother will be 4 times Jessica's age. How old is Jessica now? How old is grandmother now?
Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

Jessica is x years old
Grandmother is x+54
in one year, x+1 will be Jessica, and 4(x+1)=x+55 for grandmother
4x+4=x+55
3x=51
x=17
Jessica is 17 and her Grandmother is 71
in one year, she will be 18 and her grandmother 72, four times her age.

Answer by ikleyn(52788)

.
If Jessica is now J years old, then her grandmother, who is 54 years older, is J + 54 years old.


The condition says


(J + 54) + 1 = 4*(J + 1)


Simplify and solve for J:


J + 55 = 4J + 4

55 - 4 = 4J - J

51 = 3J  ====>  J = 51/3 = 17.


Answer.  Jessica is 17 years old.  Grandmother is 17 + 54 = 71 years old.

Solved.

---------------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.